# Can It Edit? Evaluating the Ability of Large Language Models to Follow Code Editing Instructions

### v5

@@ Abstract
Wellesley College Northeastern University
@@ 1 Introduction
Large language models of code (Code LLMs) are becoming an essential tool for software engineering practice and research. There has been significant research on synthesizing code from natural language instructions, but comparatively less attention has been given to code editing tasks. However, LLM users expect models to be capable of editing code. For example, the LMsys dataset of in-the-wild conversations with chatbots (Zheng et al., 2023) has 4,188 conversations containing code, and 831 (19%) of these involve code editing, where the user prompts the model to update code based on natural language instructions (Appendix E). In general, code editing with an LLM encompasses activities like feature addition or removal, bug fixing, and code refactoring (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023).
@@ 2 Related Work
Code Generation Benchmarks. Several benchmarks exist that test a model’s code generation ability. HumanEval and MBPP are two prominent benchmarks for evaluating LLMs in Python programming (Chen et al., 2021; Austin et al., 2021). MultiPL-E expands these benchmarks to 18+ additional programming languages (Cassano et al., 2023b). These benchmarks assess model-generated candidate completions against a series of human-authored unit tests. EvalPlus (Liu et al., 2023) utilizes mutation testing to expand the test suites of the Python benchmarks. All of these benchmarks utilize the pass@k metric, which measures the likelihood of the model generating a completion that passes all of the tests in $k$ tries; we also adopt this metric in our evaluation (§ 5.1). However, these benchmarks are limited to the evaluation of a model’s ability to generate a single function from a natural language description and do not assess code editing capabilities. HumanEvalPack (Muennighoff et al., 2023) is a benchmark designed for evaluating LLMs across various single-function code generation tasks, such as synthesis, code explanation, and bug fixing. Specifically, HumanEvalFix, a bug-fixing variant of HumanEvalPack, is extensively used for assessing the models’ capabilities in code refinement (Moon et al., 2023; Muennighoff et al., 2023). However, the instruction is fixed for every problem. SWE-Bench (Jimenez et al., 2023) evaluates LLMs across varied programming tasks including planning, retrieval, and code editing. Our work concentrates specifically on code editing tasks, aiming to more precisely guide model development. Unlike SWE-Bench, which sources its problems from GitHub PRs and issues, our benchmark is handcrafted, reducing contamination risks as seen with models like StarCoder and StarCoder2, which are trained extensively on GitHub data (Li et al., 2023b; Lozhkov et al., 2024).
@@
Code Editing Using Large Language Models. Previous studies on code editing with LLMs have predominantly focused on bug fixing (Zhang et al., 2023; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023; Joshi et al., 2023; Wei et al., 2023), a specific subset of code editing; fill-in-the-middle code completion (Bavarian et al., 2022; Fried et al., 2023; Yee & Guha, 2023; Roziere et al., 2023; Guo et al., 2024), an inference strategy that requires specific insert locations; and intrinsic code editing (Li et al., 2023a; Gupta et al., 2023), which involves editing code without a specified instruction, exerting the model’s ability to intrinsically ascertain the desired code changes. Recently, LLMs have progressed in code editing guided by natural language without specific edit locations (Hu et al., 2023; Li et al., 2023b; Muennighoff et al., 2023). However, this advancement lacks benchmark evaluations to effectively measure the models’ code editing skills. Notably, StarCoder (Li et al., 2023b), the first LLM trained on an extensive dataset of commits using the format <before><commit message><after>, we have shown enhanced code editing capabilities (§ 5). Before this study, StarCoder’s practical code editing performance had not been assessed. StarCoder2 has replaced commits with pull requests and issues, which typically include more natural language (Lozhkov et al., 2024). The recent introduction of InstructCoder (Hu et al., 2023), a model explicitly trained and evaluated for code editing, marks a significant step towards code editing with LLMs. However, its evaluation involved GPT-4-generated (OpenAI, 2023b) and human-provided labels, which raises issues regarding reproducibility and comparability in future research. Moreover, the model has not been publicly released, prohibiting us from evaluating it on our benchmark.
@@ 3 The CanItEdit Dataset
To manually construct CanItEdit, we assembled a team of eight experienced Python programmers, each with different domain expertise, and appointed one as the lead. Our objective was to fill each change category with 35 problems, with 105 problems total. Before starting, we provided the team with verbal and written guidance, a standard template, and an example problem. They were instructed to begin by writing a brief description of the problem and the applied changes for review and refinement by the lead. Next, they were tasked to write the ’before’ code segment and hidden test suite, followed by the ’after’ code segment, along with the instructions. The lead initially reviewed all problems in development and they were additionally reviewed by the entire team in weekly meetings. Upon completion of a problem, the lead generated sample completions to ensure that the failures and successes were reasonable and consistent with the problem’s intent.
The team also dedicated significant effort to developing comprehensive test suites for each problem, which incorporated a variety of testing techniques such as unit tests, property-based testing, mocking, fuzzing, and integration tests. These suites were designed to rigorously evaluate whether the ’after’ segment met the problem requirements while ensuring the ’before’ code did not. To confirm the completeness and correctness of the test suites, we created an automated verification pipeline that ensured $100\%$ line coverage and that the suite passed all tests with the ‘after’ code while failing at least one with the ‘before’ code. The team also manually reviewed the tests to ensure correctness and completeness.
@@ 4 Fine-Tuning
We describe our approach to fine-tuning Code LLMs for code editing tasks, focusing on the DeepSeekCoder-Base family (Guo et al., 2024), a variant of CodeLlama (Roziere et al., 2023) trained on 2 trillion tokens of GitHub code and natural language, using StarCoder’s filtering rules (Li et al., 2023b). These models, top-performing in code generation and open-access under a permissive license, show robust performance on CanItEdit without specific training for instructional tasks (§ 5).
@@
Originally, the dataset contained 56,025 commits, and after applying the filtering steps, we are left with 22,602. As shown by Figure 1(a) and Table 2, the mean number of lines in the ‘before’ and ‘after’ code segments is $29.2$. The mean Levenshtein distance between the ‘before’ and ‘after’ code segments is $197.1$ characters, indicating that the changes are relatively small. We also analyze the distribution of the commit message lengths, and find that the mean token count is $10.1$, which is quite low. We further analyzed the original CommitPackFT dataset to ensure that these findings weren’t artifacts of our filtering strategy: the full dataset has similar statistics.
@@
To address the limitations of EditPackFT, we developed the Commits2023FT dataset. This dataset is filtered from 416,792 Python file changes gathered from commits in permissively licensed GitHub repositories. We name the unfiltered dataset Commits2023. Our objective is to create a dataset akin to CommitPackFT, but with more recent data and a more diverse example length distribution. We employed the same filters on this dataset as used for EditPackFT, and also applied the initial filters from CommitPackFT, which includes only commits with messages that start with an imperative verb followed by at least a noun. Additionally, we only retain one file from multi-file commits to avoid exact duplicate commit messages in our dataset. After this filtering process, we obtain a dataset comprising 24,129 Python file changes. Figure 1(b) presents a broader distribution of the number of lines in the ‘before’ and ‘after’ code segments, with an average of $119.3$. We also observe a much larger change distribution, with a mean Levenshtein distance of $406.6$, signaling that this is not only a dataset with larger code segments, but also contains more varied changes. Figure 3 illustrates a sunburst plot of the most frequent initial verbs in the commit messages of Commits2023FT, along with their corresponding root nouns. This set of verbs is slightly more varied than those in EditPackFT, featuring $199$ unique verbs in comparison to $184$. Furthermore, the token count distribution of the commit messages is twice as high and much more varied than that of EditPackFT, with a mean of $23.1$ and a standard deviation of $35.2$.
@@ 5.2 Results with Existing Models
Closed source models outperform open source models. Our evaluation indicates a significant performance disparity between open and closed models. GPT-4, despite not being specifically trained on code-related tasks, surpasses DeepSeekCoder-Instruct 33b – the leading open source model – by an average of 13% in pass@1 for both descriptive and lazy tasks. DeepSeekCoder-Instruct stands out as the only open model exceeding any closed model’s performance, surpassing GPT-3.5-Turbo for descriptive prompts.
@@ 5.3 Results after Fine-Tuning on Commits
Training Dataset Metrics
Name #Tokens #Items pass@1 ExcessCode
Commits2023FT+EditPackFT 74M 46,274 43.81 0.34 $\pm$ 0.15
Commits2023FT 62M 24,129 41.88 0.33 $\pm$ 0.17
Commits2023Raw25k 62M 25,000 38.86 0.3 $\pm$ 0.12
EditPackFT 12M 22,602 41.6 0.26 $\pm$ 0.14
Table 4: Ablation results of training DeepSeekCoder-6.7b-Base on different datasets and evaluating on CanItEdit. We show the total number of tokens and items in each dataset, as well as the pass@1 and ExcessCode metrics for both the descriptive and lazy prompts aggregated across all problems. The reported sizes of the datasets are after deduplication.
@@ A.3 Results Per Change Type
Table 6: Results of OpenAI models and EditCoder on CanItEdit per change type. We report the pass@1 and ExcessCode metrics for each change type, as well as the size of the model if available. Results for lazy and descriptive prompts are aggregated across all problems. pass@1 is abbreviated to p@1.
@@ A.4 Evaluation on HumanEvalPack
DeepSeekCoder-Instruct 33b ${}^{\dagger}$ 79.2
CodeLlama-Instruct 34b ${}^{\dagger}$ 43.8
StarCoder2 15b ${}^{\dagger}$ 51.6
StarCoderBase 15b ${}^{\dagger}$ ${}^{\ddagger}$
OctoCoder 15b ${}^{\dagger}$ ${}^{\ddagger}$
CodeLlama-Instruct 13b ${}^{\dagger}$ 23.7
@@ A.6 Prompt Templates Used in Evaluation
     <user>  You are PythonEditGPT. You will be  provided the original code snippet and  an instruction that specifies  the changes you need to make.  You will produce the changed code, based  on the original code and the instruction  given. Only produce the code, do not  include any additional prose.    ## Code Before  ‘‘‘py  def add(a, b):      return a + b  ‘‘‘    ## Instruction  Add a "sub" function that subtracts two  numbers. Also write docstrings for both  functions and change a,b to x,y.  <assistant>  ## Code After  ‘‘‘py  def add(x, y):      """Adds two numbers."""      return x + y  def sub(x, y):      """Subtracts two numbers."""      return x - y  ‘‘‘  <user>  You are PythonEditGPT. You will be  provided the original code snippet and  an instruction that specifies  the changes you need to make.  You will produce the changed code, based  on the original code and the instruction  given. Only produce the code, do not  include any additional prose.    ## Code Before  ‘‘‘py  {before}  ‘‘‘  ## Instruction  {instruction}
@@
     <system>  You are PythonEditGPT. You will be  provided the original code snippet and  an instruction that specifies  the changes you need to make.  You will produce the changed code, based  on the original code and the instruction  given. Only produce the code, do not  include any additional prose.  <user>  ## Code Before  ‘‘‘py  def add(a, b):      return a + b  ‘‘‘    ## Instruction  Add a "sub" function that subtracts two  numbers. Also write docstrings for both  functions and change a,b to x,y.  <assistant>  ## Code After  ‘‘‘py  def add(x, y):      """Adds two numbers."""      return x + y  def sub(x, y):      """Subtracts two numbers."""      return x - y  ‘‘‘  <user>  ## Code Before  ‘‘‘py  {before}  ‘‘‘  ## Instruction  {instruction}
@@
     <issue_start>username_0: I have a program in Python that I’d like to change.    Here is the code for the program:  ‘‘‘py  def add(a, b):      return a + b  ‘‘‘    The change I’d like to make is:  Add a "sub" function that subtracts two numbers.  Also write docstrings for both functions and change a,b to x,y.    Please someone help me. Can you also provide the full code with the change?  <issue_comment>username_1: Sure, no problem. I will be able to help.  I am an expert in editing Python code.    Here is the full code with the change:  ‘‘‘py  def add(x, y):      \"\"\"Adds two numbers.\"\"\"      return x + y        def sub(x, y):      \"\"\"Subtracts two numbers.\"\"\"      return x - y  ‘‘‘  Upvotes: 200<issue_comment>username_0: Thank you so much!  I have another program in Python that I’d like to change.    Here is the code for the program:  ‘‘‘py  {before}  ‘‘‘    The change I’d like to make is:  {instruction}    Please someone help me. Can you also provide the full code with the change?  Upvotes: 100<issue_comment>username_1: Sure, no problem. I will be able to help.  I am an expert in editing Python code.    Here is the full code with the change:  ‘‘‘py  {after}  ‘‘‘

### v6

@@ Abstract
Wellesley College Northeastern University and Roblox
@@ 1 Introduction
Large language models of code (Code LLMs) are becoming an essential tool for software engineering practice and research. There has been significant research on synthesizing code from natural language instructions, but comparatively less attention has been given to code editing tasks. However, LLM users expect models to be capable of editing code. For example, the LMsys dataset of in-the-wild conversations with chatbots (Zheng et al., 2023) has 4,188 conversations containing code, and 831 (19%) of these involve code editing, where the user prompts the model to update code based on natural language instructions (Appendix E). In general, code editing with an LLM encompasses activities like feature addition or removal, bug fixing, and code refactoring (Zhang et al., 2023b; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023).
@@ 2 Related Work
Code Generation Benchmarks. Several benchmarks exist that test a model’s code generation ability. HumanEval and MBPP are two prominent benchmarks for evaluating LLMs in Python programming (Chen et al., 2021; Austin et al., 2021). MultiPL-E expands these benchmarks to 18+ additional programming languages (Cassano et al., 2023b). These benchmarks assess model-generated candidate completions against a series of human-authored unit tests. EvalPlus (Liu et al., 2023) utilizes mutation testing to expand the test suites of the Python benchmarks. EvoEval (Xia et al., 2024) addresses limitations in existing code generation benchmarks by evolving them into diverse domains. It reveals significant performance drops in LLMs compared to standard benchmarks, highlighting the need for more comprehensive evaluation methods. All of these benchmarks utilize the pass@k metric, which measures the likelihood of the model generating a completion that passes all of the tests in $k$ tries; we also adopt this metric in our evaluation (§ 5.1). However, these benchmarks are limited to the evaluation of a model’s ability to generate a single function from a natural language description and do not assess code editing capabilities. HumanEvalPack (Muennighoff et al., 2023) is a benchmark designed for evaluating LLMs across various single-function code generation tasks, such as synthesis, code explanation, and bug fixing. Specifically, HumanEvalFix, a bug-fixing variant of HumanEvalPack, is extensively used for assessing the models’ capabilities in code refinement (Moon et al., 2023; Muennighoff et al., 2023). However, the instruction is fixed for every problem. SWE-Bench (Jimenez et al., 2023) evaluates LLMs across varied programming tasks including planning, retrieval, and code editing. Our work concentrates specifically on code editing tasks, aiming to more precisely guide model development. Unlike SWE-Bench, which sources its problems from GitHub PRs and issues, our benchmark is handcrafted, reducing contamination risks as seen with models like StarCoder and StarCoder2, which are trained extensively on GitHub data (Li et al., 2023b; Lozhkov et al., 2024).
@@
Code Editing Using Large Language Models. Previous studies on code editing with LLMs have predominantly focused on bug fixing (Zhang et al., 2023b; Moon et al., 2023; Shinn et al., 2023; Chen et al., 2023; Olausson et al., 2023; Jin et al., 2023; Joshi et al., 2023; Wei et al., 2023; Li et al., 2022), a specific subset of code editing; fill-in-the-middle code completion (Bavarian et al., 2022; Fried et al., 2023; Yee & Guha, 2023; Roziere et al., 2023; Guo et al., 2024a; Zhang et al., 2023a), an inference strategy that requires specific insert locations; and intrinsic code editing (Li et al., 2023a; Gupta et al., 2023), which involves editing code without a specified instruction, exerting the model’s ability to intrinsically ascertain the desired code changes. Recently, LLMs have progressed in code editing guided by natural language without specific edit locations (Hu et al., 2023; Li et al., 2023b; Muennighoff et al., 2023). However, this advancement lacks benchmark evaluations to effectively measure the models’ code editing skills. Notably, StarCoder (Li et al., 2023b), the first LLM trained on an extensive dataset of commits using the format <before><commit message><after>, we have shown enhanced code editing capabilities (§ 5). Before this study, StarCoder’s practical code editing performance had not been assessed. StarCoder2 has replaced commits with pull requests and issues, which typically include more natural language (Lozhkov et al., 2024). CodeEditorBench (Guo et al., 2024b) evaluates code editing using competitive programming problems. Unlike CanItEdit, it uses fixed prompts and lacks diverse topics and dual instruction sets. The recent introduction of InstructCoder (Hu et al., 2023), a model explicitly trained and evaluated for instructional code editing, marks a significant step towards code editing with LLMs. However, its evaluation involved GPT-4-generated (OpenAI, 2023b) and human-provided labels, which raises issues regarding reproducibility and comparability in future research. Moreover, the model has not been publicly released, prohibiting us from evaluating it on our benchmark.
@@ 3 The CanItEdit Dataset
To manually construct CanItEdit, we assembled a team of eight experienced Python programmers, each with different domain expertise, and appointed one as the lead. The team, recruited through academic networks, included researchers and engineers from various fields who use Python extensively. Our objective was to fill each change category (corrective, perfective, and adaptive) with 35 problems, for a total of 105 problems. The change category is determined by the type of edit required to transform the ‘before’ code segment into an ’after’ code segment that passes the hidden test suite.
Before starting, we provided the team with verbal and written guidance, a standard template, and an example problem. They were instructed to begin by writing a brief description of the problem and the applied changes for review and refinement by the lead. Next, they were tasked to write the ’before’ code segment and hidden test suite, followed by the ’after’ code segment, along with both lazy and descriptive instructions. The lead initially reviewed all problems in development, and they were additionally reviewed by the entire team in weekly meetings. Upon completion of a problem, the lead generated sample completions to ensure that the failures and successes were reasonable and consistent with the problem’s intent.
The team dedicated significant effort to developing comprehensive test suites for each problem, incorporating a variety of testing techniques such as unit tests, property-based testing, mocking, fuzzing, and integration tests. These suites were designed to rigorously evaluate whether the ’after’ segment met the problem requirements while ensuring the ’before’ code did not. To confirm the completeness and correctness of the test suites, we created an automated verification pipeline that ensured 100% line coverage and that the suite passed all tests with the ‘after‘ code while failing at least one with the ‘before‘ code. The team also manually reviewed the tests to ensure correctness and completeness.
Annotators were encouraged to create problems of varying difficulty and were assigned specific topics and change types. The lead provided example problems and guidance on writing unambiguous instructions to ensure fair model evaluation. All instructions and several model completions were reviewed by the project lead to maintain consistency and clarity across the dataset.
@@ 4 Fine-Tuning
We describe our approach to fine-tuning Code LLMs for code editing tasks, focusing on the DeepSeekCoder-Base family (Guo et al., 2024a), a variant of CodeLlama (Roziere et al., 2023) trained on 2 trillion tokens of GitHub code and natural language, using StarCoder’s filtering rules (Li et al., 2023b). These models, top-performing in code generation and open-access under a permissive license, show robust performance on CanItEdit without specific training for instructional tasks (§ 5).
@@
Originally, the dataset contained 56,025 commits, and after applying the filtering steps, we are left with 22,602. As shown by Figure 2(a) and Table 2, the mean number of lines in the ‘before’ and ‘after’ code segments is $29.2$. The mean Levenshtein distance between the ‘before’ and ‘after’ code segments is $197.1$ characters, indicating that the changes are small relative to CanItEdit’s $302.1$ characters. We also analyze the distribution of the commit message lengths, and find that the mean token count is $10.1$, which is quite low compared to CanItEdit’s $81.7$ tokens for the descriptive and $35.6$ tokens for the lazy instructions. We further analyzed the original CommitPackFT dataset to ensure that these findings weren’t artifacts of our filtering strategy: the full dataset has similar statistics.
@@
To address the limitations of EditPackFT, we developed the Commits2023FT dataset. This dataset is filtered from 416,792 Python file changes gathered from commits in permissively licensed GitHub repositories. We name the unfiltered dataset Commits2023. Our objective is to create a dataset akin to CommitPackFT, but with more recent data and a more diverse example length distribution. We employed the same filters on this dataset as used for EditPackFT, and also applied the initial filters from CommitPackFT, which includes only commits with messages that start with an imperative verb followed by at least a noun.
@@
Additionally, we only retain one file from multi-file commits to avoid exact duplicate commit messages in our dataset. After this filtering process, we obtain a dataset comprising 24,129 Python file changes. Figure 2(b) presents a broader distribution of the number of lines in the ‘before’ and ‘after’ code segments, with an average of $119.3$. We also observe a much larger change distribution, with a mean Levenshtein distance of $406.6$, signaling that this is not only a dataset with larger code segments, but also contains more varied changes. Figure 3 illustrates a sunburst plot of the most frequent initial verbs in the commit messages of Commits2023FT, along with their corresponding root nouns. This set of verbs is slightly more varied than those in EditPackFT, featuring $199$ unique verbs in comparison to $184$. Furthermore, the token count distribution of the commit messages is twice as high and much more varied than that of EditPackFT, with a mean of $23.1$ and a standard deviation of $35.2$.
@@ 5.2 Results with Existing Models
Closed source models outperform open source models. Our evaluation indicates a significant performance disparity between open and closed models. GPT-4, despite not being primarily trained on code-related tasks but general instruction-following, surpasses DeepSeekCoder-Instruct 33b – the leading open source model – by an average of 13% in pass@1 for both descriptive and lazy tasks. DeepSeekCoder-Instruct stands out as the only open model exceeding any closed model’s performance, surpassing GPT-3.5-Turbo for descriptive prompts.
@@ 5.3 Results after Fine-Tuning on Commits
Training Dataset Descriptive Lazy
Name #Tokens #Items pass@1 EC pass@1 EC
Commits2023FT+EditPackFT 74M 46,274 48.33 0.36 39.29 0.32
Commits2023FT 62M 24,129 46.57 0.42 37.19 0.22
Commits2023Raw25k 62M 25,000 42.57 0.41 35.14 0.17
EditPackFT 12M 22,602 47.14 0.40 36.05 0.07
Table 4: Ablation results of training DeepSeekCoder-6.7b-Base on different datasets and evaluating on CanItEdit. We show the total number of tokens and items in each dataset, as well as the pass@1 and ExcessCode (EC) metrics for both the descriptive and lazy prompts. The reported sizes of the datasets are after deduplication.
@@ 6 Conclusion
Acknowledgements
This work was partially supported by the United States National Science Foundation (Awards SHF-2052696 and SES-2326174). Federico Cassano was affiliated with Roblox for a significant part of his work on this project. We thank Northeastern Research Computing and Joydeep Biswas for providing computing resources.
@@ A.3 Results Per Change Type
Table 6: Results of OpenAI models and EditCoder on CanItEdit per change type. We report the pass@1 and ExcessCode metrics for each change type, as well as the size of the model if available. Results for lazy and descriptive prompts are aggregated across all problems. pass@1 is abbreviated to p@1. Closed models are the rows highlighted in gray.
@@ A.4 Evaluation on HumanEvalPack
DeepSeekCoder-Instruct 33b 47.5 79.2
CodeLlama-Instruct 34b 36.5 43.8
StarCoder2 15b 48.6 51.6
StarCoderBase 15b 25.6 33.6
OctoCoder 15b 30.4 35.1
CodeLlama-Instruct 13b 19.4 23.7
@@ Results
Correlation with CanItEdit
We hypothesize that the performance of models on CanItEdit is correlated with their performance on general code generation tasks. Intuitively, if a model exhibits strong capabilities in code generation tasks, it is likely to perform well on code editing tasks as well. We quantify the correlation between the performance of models on CanItEdit and HumanEvalPack by calculating a Pearson correlation coefficient between the pass@1 metric of models on CanItEdit and HumanEvalPack Synthesize, utilizing models on which we have results for both benchmarks. We find a correlation of 0.661, indicating a moderate positive correlation between the two benchmarks.
@@ A.6 Prompt Templates Used in Evaluation
     <user>  You are PythonEditGPT. You will be  provided the original code snippet and  an instruction that specifies  the changes you need to make.  You will produce the changed code, based  on the original code and the instruction  given. Only produce the code, do not  include any additional prose.    ## Code Before  ```py  def add(a, b):      return a + b  ```    ## Instruction  Add a "sub" function that subtracts two  numbers. Also write docstrings for both  functions and change a,b to x,y.  <assistant>  ## Code After  ```py  def add(x, y):      """Adds two numbers."""      return x + y  def sub(x, y):      """Subtracts two numbers."""      return x - y  ```  <user>  You are PythonEditGPT. You will be  provided the original code snippet and  an instruction that specifies  the changes you need to make.  You will produce the changed code, based  on the original code and the instruction  given. Only produce the code, do not  include any additional prose.    ## Code Before  ```py  {before}  ```  ## Instruction  {instruction}
@@
     <system>  You are PythonEditGPT. You will be  provided the original code snippet and  an instruction that specifies  the changes you need to make.  You will produce the changed code, based  on the original code and the instruction  given. Only produce the code, do not  include any additional prose.  <user>  ## Code Before  ```py  def add(a, b):      return a + b  ```    ## Instruction  Add a "sub" function that subtracts two  numbers. Also write docstrings for both  functions and change a,b to x,y.  <assistant>  ## Code After  ```py  def add(x, y):      """Adds two numbers."""      return x + y  def sub(x, y):      """Subtracts two numbers."""      return x - y  ```  <user>  ## Code Before  ```py  {before}  ```  ## Instruction  {instruction}
@@
     <issue_start>username_0: I have a program in Python that I’d like to change.    Here is the code for the program:  ```py  def add(a, b):      return a + b  ```    The change I’d like to make is:  Add a "sub" function that subtracts two numbers.  Also write docstrings for both functions and change a,b to x,y.    Please someone help me. Can you also provide the full code with the change?  <issue_comment>username_1: Sure, no problem. I will be able to help.  I am an expert in editing Python code.    Here is the full code with the change:  ```py  def add(x, y):      \"\"\"Adds two numbers.\"\"\"      return x + y        def sub(x, y):      \"\"\"Subtracts two numbers.\"\"\"      return x - y  ```  Upvotes: 200<issue_comment>username_0: Thank you so much!  I have another program in Python that I’d like to change.    Here is the code for the program:  ```py  {before}  ```    The change I’d like to make is:  {instruction}    Please someone help me. Can you also provide the full code with the change?  Upvotes: 100<issue_comment>username_1: Sure, no problem. I will be able to help.  I am an expert in editing Python code.    Here is the full code with the change:  ```py  {after}  ```